OpenStudy (2yoututu11):

Which is the equation of the line? A. x – 2y = 6 B. 6x – 3y = 6 C. –3x + 6y = 6 D. –2x + y = 6

OpenStudy (anonymous):

a i think

OpenStudy (anonymous):

A

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

since the y intercept is -3

OpenStudy (anonymous):

indeed

OpenStudy (amistre64):

the intercepts are easy to spot if you put it in this form (called the intercept form) \[ax + by = c\] \[\frac{ax}{c} + \frac{by}{c} = \frac{c}{c}\] \[\frac{x}{c/a} + \frac{y}{c/b} = 1\] hence yint = c/b xint = c/a

OpenStudy (2yoututu11):